# EMGauss: Continuous Slice-to-3D Reconstruction via Dynamic Gaussian Modeling in Volume Electron Microscopy

### v1

@@ 4.2 Isotropic Reconstruction for Simulated
CycleGAN-IR[cycleganir] 22.05 0.491 0.856 22.39 0.554 0.856
EMDiffuse[emdiffuse] 23.34 0.519 0.899 24.10 0.514 0.878
@@
We also test all the methods on xz/yz-slices, as shown in Figure 5. It can be seen that, though trained on the xz/yz-slices, EMDiffuse and CycleGAN-IR still generate xz/yz-slices with significant artifacts. This indicates that the downsampled xz/yz-slices have significant differences from the original slice, thus methods based on it are not able to generate satisfying results. Instead, though our method is trained directly on xy-slices, thanks to the 3D continuous deformation field, our model is capable of modeling the continuous deformation of the volume, and therefore can generate smoother and realistic results.
@@ 4.3 Isotropic Reconstruction for Real Anisotropic
To demonstrate the generalizability on the real scenes of our method, we evaluate all the methods on the real anisotropic Volume EM dataset FANC [phelps2021reconstruction]. The results are presented in Figure 6. Compared with CycleGAN-IR and EMDiffuse, our method achieves the most accurate and detailed results, demonstrating that in a real dataset, our method can still perform effectively. Moreover, only our method is capable of generating continuously along the z-axis, which means only our method can produce images at an arbitrary time $t$. Figure 6 shows that, at an arbitrary timestamp $t=0.53$, CycleGAN-IR and EMDiffuse fail to generate results since they rely on discrete xz/yz slices. However, our method models the whole 3D volume and thus can generate at an arbitrary timestamp.
4.4 Ablation Study
We conduct an ablation study that eliminates the teacher-student module to demonstrate the necessity of it. From Table 4, the teacher-student module largely improves the generation quality, enhancing the structural details. We also test our method on different anistropy ratios, including $\times 4$ and $\times 8$. Please refer to the supplementary materials for details.
Table 4: Ablation studies on the teacher-student bootstrapping. The results are averaged over the two isotropic datasets. Note that in the “w/o teacher-student module” experiment, the model suffers from overfitting issues; thus, we select the best result across all training iterations for it.
@@ 5 Discussion
Our method provides a general and flexible framework for reconstructing 3D structures from planar-scanned 2D slices, with a particular focus on applications in vEM. Unlike prior approaches that rely on anisotropic priors or heuristic slice-to-slice matching, our formulation reconstructs the underlying 3D volume in a more principled manner by explicitly modeling a continuous deformation field directly from the high-resolution xy slices. This design not only yields smoother inter-slice transitions but also enables a more faithful recovery of morphological variations across depth. Despite its advantages, our framework still exhibits several limitations. A primary drawback is that the number of Gaussian splatting primitives can grow significantly in certain challenging regions—especially when the input slices contain substantial noise. The resulting increase in Gaussian density may lead to excessive memory consumption. A practical remedy is to incorporate a lightweight denoising module before our reconstruction pipeline, which can stabilize the optimization and prevent the uncontrolled proliferation of Gaussian primitives. In addition, future extensions may explore adaptive Gaussian pruning, multi-scale initialization strategies, or joint learning with image-space regularizers to further improve robustness and computational efficiency.
@@ 6 Conclusion
We introduced EMGauss, a simple yet powerful framework for anisotropic volume reconstruction in volume electron microscopy. By recasting slice interpolation as dynamic 3D scene rendering with deformable Gaussian splatting, EMGauss removes the strong isotropy assumptions commonly embedded in diffusion- and GAN-based approaches and provides a more principled, geometry-aware formulation of the reconstruction problem. Leveraging a self-supervised teacher–student bootstrapping scheme, our method learns directly from the target anisotropic slices without the need for external isotropic volumes or paired HR–LR data. This enables continuous, high-fidelity slice synthesis while suppressing noise and artifacts that often arise in purely 2D or patch-based formulations. The resulting reconstructions demonstrate improved structural coherence across depth and preserve fine morphological details that are critical in vEM analysis. More broadly, EMGauss establishes a general and modality-agnostic paradigm for slice-to-volume reconstruction. Its combination of deformable Gaussians, continuous depth modeling, and self-supervised learning offers a promising foundation for high-throughput 3D imaging in diverse scientific domains. We believe that the proposed framework opens up new possibilities for scalable, robust, and data-efficient 3D reconstruction pipelines across a wide range of planar acquisition systems.

### v2

@@ 4.2 Isotropic Reconstruction for Simulated
Datasets Isotropy Resolution (nm) Anisotropy ratio Issue
EPFL [22] $\checkmark$ $5\times 5\times 5$ $6\times$ brain
FIB-25 [31] $\checkmark$ $8\times 8\times 8$ $6\times$ brain
FANC [28] $\times$ $4\times 4\times 40$ $10\times$ nerve
@@
CycleGAN-IR[6] 22.05 0.491 0.856 22.39 0.554 0.856
EMDiffuse[21] 23.34 0.519 0.899 24.10 0.514 0.878
IsoVEM [13] 23.91 0.597 0.8558 21.51 0.546 0.8456
@@
We further evaluate all methods on reconstructed xz/yz slices, as shown in Figure 5. Although EMDiffuse [21] and CycleGAN-IR [6] are trained using xz/yz slices, the generated results still exhibit noticeable artifacts and structural inconsistencies. This suggests that the downsampled xz/yz slices differ substantially from the original slices, making it difficult to accurately model the underlying volumetric structure. In contrast, although our method is trained only on xy slices, the proposed 3D continuous deformation field enables the model to capture coherent volumetric deformation across the entire volume. As a result, our method produces smoother and more realistic reconstructions on the xz/yz views.
@@ 4.3 Isotropic Reconstruction for Real Anisotropic
To demonstrate the generalizability of our method to real-world scenarios, we evaluate all methods on the real anisotropic Volume EM dataset FANC [28]. The results are shown in Figure 6. Compared with CycleGAN-IR [6] and EMDiffuse [21], our method produces more accurate and structurally detailed reconstructions, indicating that it remains effective on real anisotropic data. Moreover, only our method enables continuous generation along the $z$-axis, allowing images to be synthesized at an arbitrary timestamp $t$. As illustrated in Figure 6, when evaluated at an arbitrary timestamp $t=0.53$, CycleGAN-IR [6] and EMDiffuse [21] cannot produce valid results because they rely on discrete xz/yz slice interpolation. In contrast, our method models the entire 3D volume and can therefore generate consistent slices at any timestamp.
Table 3: Video segmentation results on EPFL datasets with synthetic anisotropy ratio of $\times 6$. Segmentation is performed on the reconstructed isotropic slice sequences using SAM2 [29], and IoU is computed against the ground-truth intermediate slices.
method CycleGAN-IR EMDiffuse Ours
IoU 0.9099 0.9555 0.9687
Table 4: Ablation studies on key components of our framework. The results are averaged over the two isotropic datasets.
@@
4.4 Downstream Segmentation Results
Isotropic slice interpolation is essential for accurate 3D reconstruction in volume electron microscopy (vEM). To evaluate downstream utility, we follow [21] and perform video-based segmentation on the reconstructed volumes. Specifically, we adopt Segment Anything Model 2 (SAM2) [29] to segment cellular structures across continuous xy slices. With ground-truth intermediate slices available, we compute their segmentation masks and measure the Intersection-over-Union (IoU) between predictions and ground truth to assess segmentation accuracy. The results are summarized in Table 3. We further provide 3D reconstruction visualizations and comparisons in the supplementary video. These experiments collectively demonstrate that high-quality anisotropy correction is crucial for downstream 3D vEM analysis, and that our reconstruction pipeline yields more coherent structural continuity and higher morphological fidelity than existing methods.
4.5 Ablation Study
We conduct ablation studies to systematically evaluate the contribution of key components in our framework, including the teacher–student module, training pipeline design, and Gaussian configuration. As shown in Table 4, removing the teacher-student module leads to a noticeable degradation in generation quality and structural fidelity. Skipping the warm-up stage results in coarse initialization and compensatory Gaussian proliferation, while omitting joint training yields low-quality pseudo-labels and suboptimal convergence. For our Gaussian configuration, we adopt dynamic opacity with time-independent rotation. Static opacity introduces noise when modeling structural appearance and disappearance, whereas dynamic rotation causes temporal jitter. Further ablations on different anisotropy ratios ($\times 4$ and $\times 8$) and the number of training slices per bundle are provided in the supplementary materials.
@@ 5 Discussion
Our method provides a general and flexible framework for reconstructing 3D structures from planar-scanned 2D slices, with a particular focus on applications in vEM. Unlike prior approaches that rely on anisotropic priors or heuristic slice-to-slice matching, our formulation reconstructs the underlying 3D volume in a more principled manner by explicitly modeling a continuous deformation field directly from the high-resolution xy slices. This design not only yields smoother inter-slice transitions but also enables a more faithful recovery of morphological variations across depth. Despite its advantages, our framework still exhibits several limitations. A primary drawback is that the number of Gaussian splatting primitives can grow significantly in certain challenging regions—especially when the input slices contain substantial noise. The resulting increase in Gaussian density may lead to excessive memory consumption. A practical remedy is to incorporate a lightweight denoising module before our reconstruction pipeline, which can stabilize the optimization and prevent the uncontrolled proliferation of Gaussian primitives. In addition, future extensions may explore adaptive Gaussian pruning or joint learning with image-space regularizers to further improve robustness and computational efficiency.
@@ 6 Conclusion
We introduced EMGauss, a simple yet powerful framework for anisotropic volume reconstruction in volume electron microscopy. By recasting slice interpolation as dynamic 3D scene rendering with deformable Gaussian splatting, EMGauss removes the strong isotropy assumptions commonly embedded in previous approaches and provides a more principled, geometry-aware formulation of the reconstruction problem. Leveraging a self-supervised teacher–student bootstrapping scheme, our method learns directly from the target anisotropic slices without requiring external isotropic volumes or paired HR–LR data. This enables continuous, high-fidelity slice synthesis while suppressing artifacts that often arise in purely 2D formulations. Experiments demonstrate improved structural coherence across depth while preserving fine morphological details critical for vEM analysis. More broadly, EMGauss establishes a general and modality-agnostic paradigm for slice-to-volume reconstruction. The combination of deformable Gaussians, continuous depth modeling, and a self-supervised learning strategy offers a promising foundation for high-throughput 3D imaging in diverse scientific domains, enabling stable reconstruction without relying on large paired datasets or modality-specific supervision. We believe that the proposed framework opens up new possibilities for scalable, robust, and data-efficient 3D reconstruction pipelines across a wide range of planar acquisition systems.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant 62250062), the Smart Grid National Science and Technology Major Project (Grant 2024ZD0801200), the Shanghai Municipal Science and Technology Major Project (Grant 2021SHZDZX0102), and the Fundamental Research Funds for the Central Universities.